Find an equation of the line having the given slope and contaning the given point.
m=0,(0,-6)
I came up with this y-6=0x-0
but I know this can not be right!

Answer by nerdybill(7384)

You are very close!
.
Given:
m=0,(0,-6)
.
The "slope intercept" form of a line is:
y = mx + b
where
m = slope
b = y-intercept
.
In this problem they gave you a slope and a point -- plug it into the above formula and solve for 'b'.
.
y = mx + b
-6 = 0(0) + b
-6 = b
.
Now, we have:
m=0
b=-6
.
plug it all back into:
y = mx + b
y = 0x + (-6)
y = -6 <<--this is your equation